Anne and Nancy use a metal alloy that is 24.9% copper to make jewelry. How many ounces of a 20% alloy must be mixed with a 27% alloy to form 100 ounces of the desired alloy?
Answer by checkley77(12844)

.27x+.20(100-x)=.249*100
.27x+20-.20x=24.9
.07x=24.9-20
.07x=4.9
x=4.9/.07
x=70 oz. of 27% alloy is needed.
100-70=30 oz. of 20% alloy is needed.
proof:
.27*70+.20*30=24.9
18.9+6=24.9
24.9=24.9